A common mineral in Earth's crust has the chemical composition 18.14% Mn, 60.719%, and 21.14% O by mass. What is it's empirical formula (MnxWyOz) ?
Answer #1

You have to divide each percentage by the mass of the element

mass of Mn 54.9

mass of Oxygen 16

mass of W is 183.84

Mn; 18.14 / 54.9 = 0.33

W; 60.719 / 183.84 = 0.33

O, 21.14 / 16 = 1.32

now divide everything by 0.33 (the smallest value)

Mn, 0.33 / 0.33 = 1

W = 0.33/0.33/1

O = 1.32 / 0.33 = 4

The formula of the compound is

MnWO4.

*hope it helps =)